# Equivariant Evidential Deep Learning for Interatomic Potentials

### v1

@@ 3.3 Spectral stabilization via coefficient-space damping
While Eq. (10) guarantees geometric validity, large-magnitude $\mathcal{S}_{i}$ may lead to ill-conditioned $\mathbf{\Sigma}_{0,i}$ due to the exponential growth/shrinkage of eigenvalues. To stabilize training without explicit eigendecomposition, we damp the coefficients in representation space before constructing $\mathcal{S}_{i}$.
@@ 3.4 Training objective
The evidential model outputs a nominal covariance $\mathbf{\Sigma}_{0,i}$ and two evidence parameters $(\nu_{i},\kappa_{i})$, optimized by maximum likelihood estimation with a regularization term.
@@
Minimizing the negative log-likelihood of Eq. (7) yields the training loss in Eq. (13). For completeness, the expanded form used in the implementation is reported in Appendix C.
@@
Algorithm 1 Forward and loss on atom $i$
@@ 4 Experiments
We evaluate $\text{e}^{2}$IP across molecular dynamics benchmarks to characterize both predictive performance and the usability of its uncertainty estimates. Specifically, we study (i) probabilistic reliability and calibration on ab initio liquid water against a deep ensemble baseline, (ii) sample efficiency and geometric consistency on supramolecular MD22 systems with non-equivariant evidential baselines and an equivariance ablation, (iii) generalization under distribution shift and backbone migration on the Out of Distribution (OOD) Silica glass benchmark, and (iv) numerical robustness of full-covariance learning via spectral-stabilization ablations. Unless otherwise stated, we use AlphaNet (Yin et al., 2025) as the $SE(3)$-equivariant backbone.
@@ 4.1 Evaluation on Ab Initio Liquid Water Thermodynamics
We evaluate $\text{e}^{2}$IP on the ab initio liquid water benchmark of Cheng et al. (Cheng et al., 2019) (revPBE0-D3). Liquid water is a demanding condensed-phase setting: strong anharmonic fluctuations and nuclear quantum effects induce substantial, configuration-dependent variability in atomic forces. This benchmark stress-tests whether UQ remains reliable and actionable under realistic physical noise, i.e., whether predicted confidence levels agree with empirical frequencies which is a prerequisite for uncertainty-aware workflows such as risk control in long-timescale molecular dynamics, anomaly screening, and ranking decisions in active learning.
We use an 80%/10%/10% train/validation/test split and report mean $\pm$ std over five random seeds. We compare against an ensemble baseline consisting of five independently trained models with different random seeds, reporting both its raw output (Ensemble(raw)) and a post-hoc calibrated variant (Ensemble(cal)). The calibration fits a single isotropic scale $\sigma^{2}$ on the validation set to target $\sim 0.9$ coverage and is fixed for test evaluation (Appendix G).
@@
We report point accuracy (Force MAE), probabilistic quality (Energy Score, ES (Gneiting and Raftery, 2007); negative log-likelihood, NLL), and calibration diagnostics (Coverage@95, Calibration Error, CE, and PIT; Fig. 2) (Table 1). While ES and NLL are both proper scoring rules, NLL is more sensitive to overconfident failures and distributional mis-scaling, whereas ES emphasizes average multivariate distribution matching. PIT assesses distributional calibration: a calibrated model yields uniform PIT values, hence $\hat{F}_{U}(p)-p\approx 0$. The results show a clear accuracy–efficiency–reliability trade-off: the ensemble achieves lower MAE via model averaging but incurs substantially higher inference latency, while $\text{e}^{2}$IP attains competitive accuracy with a single forward pass and yields an approximately $4.5\times$ per-atom speedup, which is crucial for large-scale or long-timescale simulations.
@@
Ensemble(raw) 0.386 $\pm$ 0.012 0.64 $\pm$ 0.00 0.255 0.380 237.62 0.463 $\pm$ 0.067
Ensemble(cal) 0.386 $\pm$ 0.012 0.67 $\pm$ 0.00 0.931 0.256 2.74 0.463 $\pm$ 0.067
[…]
More importantly, probabilistic and calibration metrics reveal qualitative differences beyond point error. Ensemble(raw) is severely overconfident (reliability/PIT), leading to extremely low Coverage@95 and extremely large NLL, indicating that its raw uncertainty is not directly usable in practice. Ensemble(cal) reduces global overconfidence via isotropic scaling, but a single scalar cannot capture the configuration-dependent and orientation-specific error structure in condensed phases. In contrast, $\text{e}^{2}$IP predicts a full covariance tensor end-to-end, explicitly modeling anisotropic and correlated uncertainty in 3D force space; this yields markedly better probabilistic consistency (lower NLL and CE/PIT) and fewer high-confidence-but-wrong predictions, making it more suitable for uncertainty-weighted simulation. Fig. 2 further shows stronger alignment between predicted epistemic scale and true force error, directly improving ranking-based decisions such as active learning (Kulichenko et al., 2023; Jung et al., 2024).
Although $\text{e}^{2}$IP scores slightly lower on ES and Coverage@95, the high score of Ensemble(cal) depends on calibration, while the uncalibrated Ensemble(raw) only achieves 0.255 on this metric. Additionally, calibration also results in a slight increase in ES for Ensemble. Considering the Spearman correlation, CE, and NLL metrics, where $\text{e}^{2}$IP significantly outperforms ensembles, we find that $\text{e}^{2}$IP offers a more favorable accuracy–efficiency–reliability balance for uncertainty-aware MD workflows.
@@ 4.2 Sample Efficiency with Equivariance
Across both systems, $\text{e}^{2}$IP consistently exhibits improved data efficiency, with the largest and most uniform gains on DWCT. This behavior aligns with the strong geometric anisotropy and extended interactions in nanotube systems, where enforcing rotational equivariance provides a powerful inductive bias that improves generalization under limited data. Removing equivariance (Ours w/o equiv.) degrades performance despite sharing the same training objective, indicating that the gains are not driven solely by the loss design. On Buckyball-Catcher, $\text{e}^{2}$IP also improves sample efficiency, although the margin is less pronounced. We note that $\text{e}^{2}$IP maintains stable Spearman correlation as data decreases, suggesting that it preserves relative force ordering while improving accuracy. We further probe this inductive bias by evaluating on randomly rotated DWCT test configurations, where equivariance becomes essential and $\text{e}^{2}$IP achieves a larger margin.
@@
Table 2: Change in Spearman correlation under random $SO(3)$ rotations on DWCT. We report $\Delta\rho=\rho_{\text{rot}}-\rho_{\text{orig}}$ (closer to 0 is better). Best results are highlighted in bold.
@@ 4.3 Backbone-migration Evaluation on Silica Glass (OOD)
Following the dataset protocol, we subsample 1000 structures from the available training pool for training and use the remaining training structures for validation; evaluation is performed on the OOD test set. Table 3 reports force MAE together with the rank correlation between per-structure force RMSE and predicted uncertainty, where higher correlation indicates better uncertainty–error alignment for decision-making. Across both metrics, $\text{e}^{2}$IP improves over the eIP baseline, demonstrating that our uncertainty head transfers effectively to a different backbone and remains useful under a challenging distribution shift.
Table 3: Results on the Silica test set (OOD). We report force MAE and the spearman correlation coefficient between force RMSE and predicted uncertainty (higher is better). Best results are highlighted in bold.
Method Forces MAE $\downarrow$ $\rho$ $\uparrow$
eIP  (Xu et al., 2025) 1.69 0.7603
$\text{e}^{2}$IP (Ours) 1.49 0.7759
@@ Discussion.
We introduced Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a single forward pass evidential framework for uncertainty estimation in vector-valued force fields. By representing predictive uncertainty as a full $3\times 3$ SPD covariance, $\text{e}^{2}$IP captures directional structure and cross-axis correlations that scalarized formulations cannot express. More broadly, our results indicate that geometric structure plays a direct role not only in predictive accuracy, but also in making uncertainty probabilistically interpretable and actionable for downstream decision-making in molecular simulation.
A central design principle of $\text{e}^{2}$IP is to couple evidential learning with $SE(3)$-equivariant rank-2 tensor modeling, ensuring that uncertainty transforms consistently under rotations. This is realized through a Lie-algebraic covariance parameterization, in which a symmetric matrix is mapped to the SPD manifold via the matrix exponential, providing positive definiteness by construction and aligning naturally with equivariant features. Our ablations further highlight an important practical consideration: learning full covariances can lead to numerical instability under finite precision if the spectrum becomes extreme. The proposed spectral damping and evidence regularization mitigate this issue, stabilizing training and improving calibration and error–uncertainty alignment.
Finally, because the uncertainty head operates on equivariant features and outputs structured tensor components, it can be integrated with different equivariant backbones with minimal modification. This backbone-agnostic design lowers the barrier to adopting structured uncertainty modeling in existing MLIP pipelines.
@@ Appendix B Notation and validity conditions
Geometry, Lie Algebra & Optimization
$\mathcal{S}_{i}\in\mathfrak{sym}(3)$ Lie Algebra Element A symmetric matrix in the linear Lie algebra space. The network predicts this unbounded tensor to ensure unconstrained optimization.
$\exp(\mathcal{S}_{i})$ Matrix Exponential The map $\mathfrak{sym}(3)\to\mathbb{S}_{++}^{3}$ that transforms $\mathcal{S}_{i}$ into the valid SPD covariance $\Sigma_{0,i}$, guaranteeing positive definiteness and equivariance by construction.
@@ Appendix D Irrep-to-Cartesian map and the constant basis matrix $\mathbf{Q}$
This appendix specifies the implementation-dependent details omitted in Sec. 3.2. Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the coefficients in $(0e\oplus 2e)$. We map $\mathbf{z}_{i}$ to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathfrak{sym}(3)$ using a fixed Clebsch–Gordan basis.
@@
$\mathrm{vec}(\mathbf{S}_{i})=\mathbf{Q}^{\top}\mathbf{z}_{i}^{\top},\qquad\mathbf{Q}\in\mathbb{R}^{9\times 6},$ (26)
@@
$\mathcal{S}_{i}\triangleq\tfrac{1}{2}(\mathbf{S}_{i}+\mathbf{S}_{i}^{\top})\in\mathfrak{sym}(3).$ (27)

### v2

@@ 3.3 Spectral stabilization via coefficient-space damping
The construction above guarantees that $\mathbf{\Sigma}_{0,i}$ is always SPD and equivariant. However, in practice the matrix exponential amplifies extreme eigenvalues: if the network predicts large-magnitude $\mathcal{S}_{i}$, the resulting $\mathbf{\Sigma}_{0,i}$ becomes severely ill-conditioned (eigenvalue ratios $10^{5}$), causing numerical failures in the Cholesky factorization required by the NLL loss (Fig. 4, right). To prevent this without breaking equivariance or requiring explicit eigendecomposition, we damp the irrep coefficients before constructing $\mathcal{S}_{i}$.
@@ 3.4 Training objective
We now describe how the full pipeline — backbone $\to$ evidential heads $\to$ tangent-space construction $\to$ spectral damping $\to$ matrix exponential — is trained end-to-end. Algorithm 1 summarizes the forward pass. Given labeled data $\{(\mathcal{G}_{j},\{\vec{y}_{i}^{j}\})\}$ where $\mathcal{G}_{j}$ is an atomic graph and $\vec{y}_{i}^{j}\in\mathbb{R}^{3}$ is the DFT force on atom $i$, the network outputs $(\vec{\gamma}_{i},\mathbf{\Sigma}_{0,i},\nu_{i},\kappa_{i})$ per atom. The loss has two components.
@@
Minimizing the negative log-likelihood of Eq. (7) yields the training loss in Eq. (13). Intuitively, the $\log|\mathbf{\Sigma}_{0,i}|$ term penalizes overconfident covariance (too small $\mathbf{\Sigma}_{0,i}$ is costly), while the Mahalanobis term M penalizes force prediction errors weighted by the predicted uncertainty. For completeness, the expanded form used in the implementation is reported in Appendix C.
@@
Algorithm 1 Forward pass and loss computation for atom $i$. Steps 1–3 use the shared backbone; steps 4–9 are the evidential uncertainty head; steps 10–14 compute the loss.
@@ 4 Experiments
We evaluate $\text{e}^{2}$IP across five experimental settings designed to probe different aspects of uncertainty quality:
  • Sec. 4.1: Probabilistic reliability on ab initio liquid water, comparing against deep ensembles (MAP and heteroskedastic), shallow ensembles, and eIP.
  • Sec. 4.2: Sample efficiency and rotation consistency on MD22 supramolecular systems.
  • Sec. 4.3: Backbone migration and OOD generalization on the Silica glass benchmark.
  • Sec. 4.4: Standard force-field evaluation on rMD17 molecular benchmarks.
  • Sec. 4.5: Ablation of spectral stabilization.
Unless otherwise stated, we use AlphaNet (Yin et al., 2025) as the $SE(3)$-equivariant backbone. Evaluation metrics (ES, CE, NLL, Coverage, Spearman) are defined in Appendix F.
@@ 4.1 Evaluation on Ab Initio Liquid Water Thermodynamics
We evaluate on the ab initio liquid water benchmark (Cheng et al., 2019) (revPBE0-D3), a demanding condensed-phase setting where strong anharmonic fluctuations produce configuration-dependent force variability. This stress-tests whether UQ remains reliable under realistic physical noise — a prerequisite for uncertainty-aware workflows such as active learning and risk control.
We use an 80%/10%/10% train/validation/test split and report mean $\pm$ std over five random seeds. We compare against several baselines: (1) a deep ensemble (DE) consisting of five independently MAP-trained models, reported both raw (DE(raw)) and with post-hoc isotropic calibration (DE(cal), $\sigma^{2}$ fitted on validation set; Appendix G); (2) a heteroskedastic deep ensemble (MVE-DE) where each of K=5 members predicts both force mean and per-component diagonal variance, trained with Gaussian NLL loss following the standard MVE formulation (Scalia et al., 2020; Kellner and Ceriotti, 2024); (3) a shallow ensemble (SE) with shared backbone and K=5 independent output heads (Kellner and Ceriotti, 2024); and (4) the diagonal evidential baseline eIP (Xu et al., 2025).
@@
Table 1 reports point accuracy (Force MAE), probabilistic quality (Energy Score ES (Gneiting and Raftery, 2007), NLL), and calibration diagnostics (Coverage@95, Calibration Error CE, PIT; Fig. 2). The results show a clear accuracy–efficiency–reliability trade-off: DE achieves the lowest MAE via model averaging but incurs ${\sim}4.5\times$ higher inference cost, while $\text{e}^{2}$IP attains competitive accuracy in a single forward pass.
@@
DE(raw) 0.386 $\pm$ 0.012 0.64 $\pm$ 0.00 0.255 0.380 237.62 0.463 $\pm$ 0.067
DE(cal) 0.386 $\pm$ 0.012 0.67 $\pm$ 0.00 0.931 0.256 2.74 0.463 $\pm$ 0.067
MVE-DE 0.492 0.73 0.970 0.139 1.59 0.463 $\pm$ 0.067
SE(cal) 0.517 $\pm$ 0.012 0.93 $\pm$ 0.06 0.790 $\pm$ 0.071 0.134 $\pm$ 0.028 15.2 $\pm$ 5.6 0.234 $\pm$ 0.004
eIP (Xu et al., 2025) 0.499 $\pm$ 0.016 2.15 $\pm$ 2.38 0.652 $\pm$ 0.008 0.314 $\pm$ 0.008 3.13 $\pm$ 0.19 0.094 $\pm$ 0.003
$\text{e}^{2}$IP (Ours) 0.471 $\pm$ 0.010 0.68 $\pm$ 0.02 0.910 $\pm$ 0.024 0.114 $\pm$ 0.028 1.31 $\pm$ 0.03 0.102 $\pm$ 0.016
More importantly, probabilistic and calibration metrics reveal qualitative differences beyond point error. DE(raw) is severely overconfident (reliability/PIT), leading to extremely low Coverage@95 and extremely large NLL, indicating that its raw uncertainty is not directly usable in practice. DE(cal) reduces global overconfidence via isotropic scaling, but a single scalar cannot capture the configuration-dependent and orientation-specific error structure in condensed phases. In contrast, $\text{e}^{2}$IP predicts a full covariance tensor end-to-end, explicitly modeling anisotropic and correlated uncertainty in 3D force space; this yields markedly better probabilistic consistency (lower NLL and CE/PIT) and fewer high-confidence-but-wrong predictions, making it more suitable for uncertainty-weighted simulation. Fig. 2 further shows stronger alignment between predicted epistemic scale and true force error, directly improving ranking-based decisions such as active learning (Kulichenko et al., 2023; Jung et al., 2024).
To address whether $\text{e}^{2}$IP’s advantage stems from modeling input-dependent aleatoric uncertainty, we additionally compare against MVE-DE, a heteroskedastic deep ensemble where each member outputs both force mean and diagonal variance. Despite having the capacity to model aleatoric uncertainty, MVE-DE achieves worse MAE (0.492) than $\text{e}^{2}$IP (0.471) and MAP-DE (0.386), and does not outperform $\text{e}^{2}$IP on any UQ metric (ES 0.73 vs 0.68; CE 0.139 vs 0.114; NLL 1.59 vs 1.31). The shallow ensemble (SE) performs even worse (NLL 15.2), suggesting that output-layer diversification alone is insufficient for reliable multivariate force UQ.
We note that NLL uses different distributional forms: Gaussian for DE/MVE-DE/SE, and Student-$t$ for $\text{e}^{2}$IP (the native predictive from NIW marginalization). The Energy Score (ES) is a strictly proper scoring rule that depends only on samples and requires no parametric assumptions (Gneiting and Raftery, 2007), making it directly comparable across distributional families. $\text{e}^{2}$IP achieves the best ES across all methods.
Overall, combining stronger probabilistic reliability (NLL, CE, ES), better error–uncertainty ranking alignment, and substantially lower inference cost (${\sim}4.5\times$ speedup over DE), $\text{e}^{2}$IP provides more actionable uncertainty while retaining competitive accuracy on ab initio liquid water.
@@ 4.2 Sample Efficiency with Equivariance
Across both systems, $\text{e}^{2}$IP consistently exhibits improved data efficiency, with the largest and most uniform gains on DWCT. This behavior aligns with the strong geometric anisotropy and extended interactions in nanotube systems, where enforcing rotational equivariance provides a powerful inductive bias that improves generalization under limited data. Removing equivariance (Ours w/o equiv.) degrades performance despite sharing the same training objective, indicating that the gains are not driven solely by the loss design. On Buckyball-Catcher, $\text{e}^{2}$IP also improves sample efficiency, although the margin is less pronounced. We note that the non-equivariant variant can match or exceed $\text{e}^{2}$IP on raw accuracy at high data fractions on some systems; understanding when and why this occurs is an interesting open question. Nevertheless, $\text{e}^{2}$IP maintains stable Spearman correlation as data decreases, suggesting that it preserves relative force ordering while improving accuracy. Importantly, the primary advantage of equivariance is not raw MAE dominance at every data point, but rotation-consistent uncertainty: we further probe this by evaluating on randomly rotated DWCT test configurations (Table 2), where $\text{e}^{2}$IP achieves substantially smaller $|\Delta\rho|$.
@@
Table 2: Change in Spearman correlation under random $SO(3)$ rotations on DWCT. We report $|\Delta\rho|=|\rho_{\text{rot}}-\rho_{\text{orig}}|$ (closer to 0 is better). For multi-seed methods we report mean $\pm$ std. Best results are highlighted in bold.
@@ 4.3 Backbone-migration Evaluation on Silica Glass (OOD)
Following the dataset protocol, we subsample 1000 structures from the available training pool for training and use the remaining training structures for validation; evaluation is performed on the OOD test set. Table 3 reports force MAE, rank correlation, and inference time. We additionally trained a DE baseline (K=5 PaiNN models) and re-ran eIP (5 seeds) and $\text{e}^{2}$IP (4 seeds) with error bars. DE achieves the lowest MAE at $4.6\times$ inference cost (0.242 vs 0.053 ms/atom). $\text{e}^{2}$IP outperforms eIP on both MAE and Spearman, and achieves comparable Spearman to DE (0.775 vs 0.790) while maintaining single-model efficiency. This demonstrates that our uncertainty head transfers effectively to a different backbone and remains useful under a challenging distribution shift.
Table 3: Results on the Silica test set (OOD). Force MAE, Spearman $\rho$, and inference time. Mean $\pm$ std over seeds. Best in bold.
[…]
4.4 Evaluation on rMD17 Molecular Force Fields
To evaluate on standard molecular force-field benchmarks with diverse molecular sizes, we compare DE, eIP, $\text{e}^{2}$IP, and MVE-DE on two molecules from the revised MD17 dataset (Christensen and von Lilienfeld, 2020): aspirin (21 atoms) and azobenzene (24 atoms). All methods use the same AlphaNet backbone and hyperparameters. Results are reported over multiple seeds (Table 4).
Table 4: Results on rMD17. MAE, ES, CE, NLL. Mean $\pm$ std. Best in bold.
[…]
Asp. DE(cal) .187 .703 .423 3.30
MVE-DE .496 .830 .444 3.35
[…]
Azo. DE(cal) .139 .341 .327 2.34
MVE-DE .640 1.238 .472 4.93
[…]
On both molecules, $\text{e}^{2}$IP achieves the best CE, ES, and NLL. On azobenzene, $\text{e}^{2}$IP also achieves the lowest MAE (0.128 vs DE’s 0.139), demonstrating that the evidential framework can compete with ensembles on point accuracy. MVE-DE performs substantially worse than both MAP-DE and $\text{e}^{2}$IP across all metrics, with significantly degraded MAE (0.496 on aspirin, 0.640 on azobenzene) and poor calibration (CE >> 0.44). Furthermore, some MVE-DE seeds failed to converge during training (aspirin seed 1, azobenzene seeds 3–4), indicating training instability when applying heteroskedastic Gaussian NLL to autograd-derived forces — a known challenge in MLIP UQ. These results, combined with the liquid water experiments, consistently demonstrate that $\text{e}^{2}$IP’s NIW evidential framework provides a more effective joint force–uncertainty parameterization than heteroskedastic ensembles.
@@ Discussion.
$\text{e}^{2}$IP couples evidential learning with $SE(3)$-equivariant rank-2 tensor modeling: a tangent-space parameterization maps unconstrained symmetric matrices to SPD covariances via the Riemannian exponential, guaranteeing positive definiteness and rotation consistency by construction. The result is a full $3\times 3$ uncertainty tensor that captures directional structure inaccessible to diagonal formulations. Our ablations show that spectral damping is essential for stable full-covariance learning (Fig. 4), and our comparison with heteroskedastic deep ensembles (MVE-DE) demonstrates that the NIW framework achieves superior calibration without the training instability observed in Gaussian NLL-based approaches. Because the uncertainty head operates on equivariant features and outputs structured tensor components, it integrates with different equivariant backbones (AlphaNet, PaiNN) with minimal modification.
@@ Appendix B Notation and validity conditions
Geometry, Tangent Space & Optimization
$\mathcal{S}_{i}\in\mathrm{sym}(3)$ Tangent Vector A symmetric matrix in the tangent space of $\mathbb{S}_{++}^{3}$ at the identity. The network predicts this unbounded tensor to ensure unconstrained optimization.
$\exp(\mathcal{S}_{i})$ Riemannian Exponential Map The map $\mathrm{sym}(3)\to\mathbb{S}_{++}^{3}$ that transforms $\mathcal{S}_{i}$ into the valid SPD covariance $\Sigma_{0,i}$, guaranteeing positive definiteness and equivariance by construction (Moakher, 2005; Arsigny et al., 2007).
@@ Appendix D Irrep-to-Cartesian map and the constant basis matrix $\mathbf{Q}$
This appendix specifies the implementation-dependent details omitted in Sec. 3.2. Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the coefficients in $(0e\oplus 2e)$. We map $\mathbf{z}_{i}$ to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathrm{sym}(3)$ using a fixed Clebsch–Gordan basis.
@@
$\mathrm{vec}(\mathbf{S}_{i})=\mathbf{Q}^{\top}\mathbf{z}_{i},\qquad\mathbf{Q}\in\mathbb{R}^{9\times 6},$ (26)
@@
$\mathcal{S}_{i}\triangleq\tfrac{1}{2}(\mathbf{S}_{i}+\mathbf{S}_{i}^{\top})\in\mathrm{sym}(3).$ (27)